# OrchestraLLM: Efficient Orchestration of Language Models for Dialogue State Tracking

### v2

@@ Abstract
Large language models (LLMs) have revolutionized the landscape of Natural Language Processing, but are computationally expensive. To reduce the cost without sacrificing performance, previous studies have explored various approaches to harness the potential of Small Language Models (SLMs) as cost-effective alternatives to their larger counterparts. Driven by findings that SLMs and LLMs exhibit complementary strengths in a structured knowledge extraction task, this work presents a novel SLM/LLM routing framework designed to improve computational efficiency and enhance task performance. In dialogue state tracking tasks, the proposed routing framework enhances performance substantially compared to relying solely on LLMs, while reducing the computational costs by over 50%.
@@ 1 Introduction
Large Language Models (LLMs) have become versatile tools capable of tackling a wide range of tasks with only a few training examples. However, their expanding sizes have brought escalating computational demands. In contrast, more efficient Small Language Models (SLMs) often require a substantial amount of fine-tuning data to become truly effective. This work addresses scenarios where only limited task-specific data is available, making fine-tuned SLMs less dependable. Our objective is to develop a routing framework that orchestrates SLMs and LLMs, enhancing task performance while reducing computational costs.
@@ 4.2 Experimental Setting
For Prompt-DST, we use T5-base and T5-large as the backbone model for MWOZ and SGD respectively, as the latter is more complex in terms of schema and more dialogue turns. For IC-DST, we use ChatGPT as the backbone model [^2] Accessed: August–October 2023, Version: gpt-3.5-turbo-0301. with 10 in-context exemplars. We initialize the routing retriever from SenBERT (all-mpnet-base-v2). We run inference on 100 dialogues randomly sampled from validation sets of MWOZ and SGD as the held-out sets. The same 100 dialogues are used to train the retriever. For all experiments, $l=25$ is used for the positive and negative examples for contrastive learning. During inference, we randomly sample 100 turns from the held-out sets to serve as SLM pool and LLM pool respectively for MWOZ experiments and 300 turns for SGD experiments. We use $k=10$ for the majority vote and break the tie by favoring SLM.
@@ 4.4 Baselines
We compare our routing framework with existing classification-based approaches to model switching, such as those proposed by Šakota et al. (2023) and Kag et al. (2022). These existing approaches typically train a binary classifier to serve as the router. We train BERT Devlin et al. (2019) (bert-base-cased) with the expert labels in the hold-out set of dialogues with a binary classification objective to do routing as a baseline.
Cascade-Based Routing
Cascade-based approaches Chen et al. (2023); Madaan et al. (2023) typically query a SLM and redirect the instance to a LLM if the small language model is not confident enough. We choose to utilize the normalized sequence level probability of SLM output as the confidence measure. We tune the probability threshold on the hold-out-set and use the threshold to determine whether to redirect the instance to LLM during inference.
@@ 5 Analysis
5.1 Cross-Domain Generalization
Out-of-Domain (OOD) in SGD To assess the effectiveness of OrchestraLLM in generalizing to unseen domains, we present breakdown results on SGD in Figure 2. First, we observe that Prompt-DST performs better than IC-DST on in-domain dialogues but lags behind IC-DST on all other types of dialogues. This suggests that the generalization ability of Large Language Models (LLMs) is superior to Small Language Models (SLMs). All variants of OrchestraLLM outperform IC-DST in OOD scenarios, demonstrating the router’s capability to effectively dispatch instances even when they are out of the domain.
Cross-Dataset Retriever We further evaluate our proposed framework in a more challenging scenario where the router and backbone models are trained in different datasets. We train the retriever model on MWOZ holdout set dialogues and evaluate the framework on SGD testing dialogues and vice versa. Notably, our routing framework can still effectively orchestrate two LLMs with a retriever trained with a different dataset and outperforms IC-DST while also achieving computational cost savings of approximately 54% on MultiWOZ and 43% on SGD (only slightly less savings than the domain-matched model).
@@ 5.1 Cross-Domain Generalization
Test turn
[system] Do you have a cuisine or price range in mind? [user] Yes, something in the affordable price range. Also, do any of them serve Singaporean food ?
@@
Test turn
[system] Would you like to book a room? [user] That would be a massive help if you can do that for me! It’s me and my mum and we’ll be there for 2 nights.
@@
Test turn
[system] Can I be of any further assistance today? [user] Yes, I am also looking for a 3-star hotel located in the same area as the restaurant.

### v3

@@ Abstract
Large language models (LLMs) have revolutionized the landscape of Natural Language Processing, but are computationally expensive. To reduce the cost without sacrificing performance, previous studies have explored various approaches to harness the potential of Smaller Language Models (SLMs) as cost-effective alternatives to their larger counterparts. Driven by findings that SLMs and LLMs exhibit complementary strengths in a structured knowledge extraction task, this work presents a novel SLM/LLM routing framework designed to improve computational efficiency and enhance task performance. In dialogue state tracking tasks, the proposed routing framework enhances performance substantially compared to relying solely on LLMs, while reducing the computational costs by over 50%.
@@ 1 Introduction
Large Language Models (LLMs) have become versatile tools capable of tackling a wide range of tasks with only a few training examples. However, their expanding sizes have brought escalating computational demands. In contrast, more efficient Smaller Language Models (SLMs) often require a substantial amount of fine-tuning data to become truly effective. This work addresses scenarios where only limited task-specific data is available, making fine-tuned SLMs less dependable. Our objective is to develop a routing framework that orchestrates SLMs and LLMs, enhancing task performance while reducing computational costs.
@@ 4.2 Experimental Setting
For Prompt-DST, we use T5-base and T5-large as the backbone model for MWOZ and SGD respectively, as the latter is more complex in terms of schema and more dialogue turns. For IC-DST, we use ChatGPT as the backbone model [^2] Accessed: August–October 2023, Version: gpt-3.5-turbo-0301. with 10 in-context exemplars. We initialize the routing retriever from SenBERT (all-mpnet-base-v2). We run inference on 100 dialogues randomly sampled from validation sets of MWOZ and SGD as the held-out sets. The same 100 dialogues are used to train the retriever. For all experiments, $l=25$ is used for the positive and negative examples for contrastive learning. During inference, we randomly sample 100 turns from the held-out sets to serve as SLM pool and LLM pool respectively for MWOZ experiments and 300 turns for SGD experiments. [^3] We also experimented with 50 turns for MWOZ and observed less than 1% accuracy degradation. We leave it for future work to create novel strategies for expert pool instance selection. We use $k=10$ for the majority vote and break the tie by favoring SLM.
@@ 4.4 Baselines
We compare our routing framework with existing classification-based approaches to model switching, such as those proposed by Šakota et al. (2023) and Kag et al. (2022). These existing approaches typically train a binary classifier to serve as the router. We train BERT Devlin et al. (2019) (bert-base-cased) with the expert labels in the hold-out set of dialogues with a binary classification objective to do routing as a baseline. Cascade-Based Routing
Cascade-based approaches Chen et al. (2023); Madaan et al. (2023) typically query a SLM and redirect the instance to a LLM if the smaller language model is not confident enough. We choose to utilize the normalized sequence level probability of SLM output as the confidence measure. We tune the probability threshold on the hold-out-set and use the threshold to determine whether to redirect the instance to LLM during inference.
@@ 4.5.1 MultiWOZ
Router SGD MWOZ
Assignment Ratio TLB JGA Assignment Ratio TLB JGA
SenBERT 50% 65.97 60% 80.74
Task-Aware (SGD) 55% 67.25 54% 80.75
Task-Aware (MWOZ) 43% 66.57 55% 82.43
Table 4: Cross-dataset routing results of OrchestraLLM on SGD and MWOZ. We denote the % of testing turns routed to Prompt-DST (SLM) as Assignment Ratio.
@@ 5 Analysis
5.1 Cross-Domain Generalization
Out-of-Domain (OOD) in SGD To assess the effectiveness of OrchestraLLM in generalizing to unseen domains, we present breakdown results on SGD in Figure 2. First, we observe that Prompt-DST performs better than IC-DST on in-domain dialogues but lags behind IC-DST on all other types of dialogues. This suggests that the generalization ability of Large Language Models (LLMs) is superior to Smaller Language Models (SLMs). All variants of OrchestraLLM outperform IC-DST in OOD scenarios, demonstrating the router’s capability to effectively dispatch instances even when they are out of the domain.
Cross-Dataset Retriever We further evaluate our proposed framework in a more challenging scenario where the router and backbone models are trained in different datasets. We train the retriever model on MWOZ holdout set dialogues and evaluate the framework on SGD testing dialogues and vice versa. The results are shown in Table 4. Notably, our routing framework can still effectively orchestrate two LLMs with a retriever trained with a different dataset and outperforms IC-DST while also achieving computational cost savings of approximately 54% on MultiWOZ and 43% on SGD. Compared with the dataset-matched finetuning, the mismatched finetuning with a different dataset only slightly hurts the accuracy and efficiency.
@@ 5.1 Cross-Domain Generalization
Test turn [system] Do you have a cuisine or price range in mind? [user] Yes, something in the affordable price range. Also, do any of them serve Singaporean food ?
@@
Test turn [system] Would you like to book a room? [user] That would be a massive help if you can do that for me! It’s me and my mum and we’ll be there for 2 nights.
@@
Test turn [system] Can I be of any further assistance today? [user] Yes, I am also looking for a 3-star hotel located in the same area as the restaurant.
@@ 5.2 Specialty of SLM and LLM
Models Assignment TLB
T5-base N/A 73.43
T5-3B N/A 78.77
OrchestraLLM (T5-base+T5-3B) 61% 81.09
Table 6: Routing results between T5-base and T5-3B(a new LM) using an off-the-shelf SenBERT. We denote the % of testing turns routed to T5-base as Assignment.
5.3 Routing a New LM
To demonstrate the flexibility of OrchestraLLM, we provide routing results of when integrating a new LM, specifically T5-3b finetuned with Prompt-DST method on MultiWOZ 5% training data. We apply OrchestraLLM with an off-the-shelf SenBERT to route between T5-base and T5-3b. The results displayed in Table 6 underscore the adaptability of OrchestraLLM in effectively routing examples with a newly introduced LM, all without requiring any additional training of the retriever.